the local orchard has just opened a new plot of land for strawberry picking that is 500 square feet. If the length is 20 feet longer than the width, what are the dimensions of the strawberry patch?
Answer by mananth(16946)

let the width be x
The length will be x+20
The area = 500 sq.ft
..
area = L*W
x(x+20)=500
x^2+20x=500
x^2+20x-500=0
..
x1 , x2 the roots have to be found out.
x1=-20+sqrt(20^2+4*1*500) /2
x1=14.49
x2=-40-sqrt(40^2+4*1*500) /2
=--34.49
..
width =14.49 feet
length = 500/14.49
=34.5 feet